# Constrained Contextual Online Decision Making: A Unified Framework

### v1

@@ Abstract
Contextual online decision-making problems with constraints appear in a wide range of real-world applications, such as personalized recommendation with resource limits, adaptive experimental design, and decision-making under safety or fairness requirements. In this paper, we investigate a general formulation of sequential decision-making with stage-wise feasibility constraints, where at each round, the learner must select an action based on observed context while ensuring that a problem-specific feasibility criterion is satisfied. We propose a unified algorithmic framework that captures many existing constrained learning problems, including constrained bandits, active learning with label budgets, online hypothesis testing with Type I error control, and model calibration. Central to our approach is the concept of upper counterfactual confidence bounds, which enables the design of practically efficient online algorithms with strong theoretical guarantees using any offline conditional density estimation oracle. Technically, to handle feasibility constraints in complex environments, we introduce a generalized notion of the eluder dimension, extending it from the classical setting based on square loss to a broader class of metric-like probability divergences. This allows us to capture the complexity of various density function classes and characterize the utility regret incurred due to feasibility constraint uncertainty. Our result offers a principled foundation for constrained sequential decision-making in both theory and practice.
@@ 1 Introduction
Contextual sequential online experimentation plays a critical role in various domains, including online platforms (Ban et al., 2024), healthcare systems (Zhou et al., 2023), and other industries. In many real-world applications, the decision maker (DM) often operates under some constraints that must be satisfied (Min and Azizan, 2024). For example, there may be safety constraints in power systems (Ceusters et al., 2023), fairness constraints on online platforms (Verma et al., 2024), or risk constraints in portfolio management (Galichet et al., 2013). The presence of such constraints significantly complicates the sequential decision-making process, making it more challenging to manage the inherent exploration-exploitation trade-off.
To address this challenge, for different statistical tasks, a variety of adaptive constrained decision-making algorithms have been proposed, including constrained bandits (Jagerman et al., 2020), active learning (Awasthi et al., 2021), and online hypothesis testing (Malek et al., 2017). Despite the diversity of these applications, they often share a common structural pattern. In each round, the DM observes a context $x_{t}$, such as the symptoms of the patient in a healthcare setting. Based on this information, the DM selects an action $a_{t}$ that is deemed feasible with respect to some estimated constraint. Subsequently, the DM receives feedback in the form of data sampled from the underlying distributions associated with the context-action pair $(x_{t},a_{t})$ and proceeds to the next round.
@@
In this paper, we introduce a unified framework for general constrained sequential decision-making problems and present an efficient algorithm to handle it. To be specific, we summarize our contributions as follows:
  • A New General Framework: We propose a new framework to handle general online decision-making problems with stage-wise constraints. Beyond the unconstrained model  (Hu et al., 2025), our framework incorporates many sequential decision-making applications, such as constrained bandits, active learning, risk-aware bandits, and online hypothesis testing. All these problems could be addressed by replacing the abstract functionals defined in  Section 3 with specific task-driven functionals.
  • A Novel Conceptual Generalization of Eluder Dimension: Eluder dimension  (Russo and Van Roy, 2013) is a complexity measure often used to characterize the difficulty of optimal exploration in bandits. Moreover, in constrained online learning (Sridharan and Yoo, 2024), it is also used for characterizing the regret caused by conservative exploration. We extend the idea and concept of the eluder dimension  Russo and Van Roy (2013) from the reward function class with squared error to any density model class with any metric-like probability divergence; see  Definition 5.1. This new measure helps us balance aggressive and conservative exploration to handle feasibility.
  • An Efficient Unified Algorithm: We propose a unified algorithm that could easily transfer any offline density estimation oracle into a efficient sequential decision-making algorithm. Unlike in  Hu et al. (2025) where the underlying density class is restricted to a linear model, our algorithm could handle any density model class, including linear and non-linear model classes. This greatly enriches the range of problems that we can solve.
@@ 2 Related Work
We review related work from both theoretical and application perspectives, presenting them separately.
From a theoretical perspective, our work is closely related to several research areas, including general sequential decision-making, eluder dimension analysis, and online learning under constraints.
Sequential decision-making is a fundamental problem in modern machine learning. Early work such as Bastani and Bayati (2020) focused on sequential decision-making with finite high-dimensional covariates, using contextual bandits as a representative example. From a more general perspective, Foster et al. (2021) and Zhong et al. (2022) investigated the learnability of general model classes by introducing the decision-estimation coefficient and its generalizations to characterize the complexity of interactive decision-making. However, these studies focus primarily on maximizing expected rewards without addressing additional structural constraints.
In contrast, our paper provides a more general framework that accommodates sequential decision-making under stage-wise constraints. Closely related, Hu et al. (2025) proposed an algorithm for unconstrained online decision-making problems based on linear functional regression. Our framework goes further by supporting both linear and nonlinear density classes while also incorporating constraints into the decision process.
@@
The study of online learning with constraints was initiated by Mannor et al. (2009), where the learner aims to maximize expected rewards subject to long-term sample path constraints. Later on, Mahdavi et al. (2012); Castiglioni et al. (2022); Liakopoulos et al. (2019); Yu et al. (2017) studied the regret upper bound and constraint violation in online optimization with long-term constraints. Instead of studying constraint violation, our paper gives a feasibility guarantee such that with high probability, it holds for all rounds. On the other hand, Sridharan and Yoo (2024); Hutchinson et al. (2024) investigated online optimization with per-round constraints. The benchmark in these papers is the best fixed action in hindsight, which is non-contextual and the regret is static, whereas in our paper, we studied the per-context optimal decision-making problem and our benchmark is contextualized dynamic regret. In brief, we have a high probability feasibility guarantee with dynamic regret guarantee in our paper.
@@
Our work is also closely related to the body of research on online learning with general function approximation. Xu and Zeevi (2020) proposed the Upper Counterfactual Confidence Bound (UCCB) algorithm, which constructs upper confidence bounds in the policy space under general reward function approximation. Foster and Rakhlin (2020) addressed the contextual bandit problem using inverse-gap weighting policies, assuming access to an online regression oracle. More recently, Pacchiano (2024) and Jia et al. (2024) examined general contextual bandits with small noise levels and established connections between reward variance and regret. In the context of reinforcement learning, Wang et al. (2020a, b) studied regret minimization under general value function approximation. In contrast, our work focuses on general sequential decision-making with feasibility constraints, leveraging density function approximation oracles.
@@
Another related literature subfield is reinforcement learning under safety constraints. For comprehensive overviews, we refer readers to the surveys by Gu et al. (2022) and Wachi et al. (2024). Wachi and Sui (2020) investigated reinforcement learning with explicit state-wise safety constraints, and Li and Azizan (2024) studied its multi-agent variant and convergence to Nash equilibrium. Furthermore, HasanzadeZonuzy et al. (2021); Amani et al. (2021a, b) analyzed the sample complexity and regret of learning in value function constrained Markov Decision Processes (MDPs). Additionally, Miryoosefi et al. (2019) studied constrained MDPs where policies are required to lie within a convex policy set. Although these works primarily focus on value-based or policy-space constraints under tabular or linear MDPs, our framework is more general and can accommodate a much broader class of feasibility constraints and non-linear model classes.
@@
Furthermore, we remark on some connections and differences between our paper and control theory. Regret Optimal Control (Sabag et al., 2023; Goel and Hassibi, 2021b) is about minimizing the total cost relative to a clairvoyant optimal policy that has foreknowledge of the system dynamics. Recently, Martin et al. (2022) studied this problem under a safety constraint. In the literature on regret-optimal control, the primary focus has been on identifying the optimal controller and characterizing its structural properties (Sabag et al., 2021; Goel and Hassibi, 2021a; Martin et al., 2024). Moreover, optimal control focuses on minimizing cumulative costs generated by a dynamic system along a single trajectory (Li et al., 2019). In our framework, however, contexts do not have system dynamics, so we aim to maximize cumulative rewards through repeated online experimentation, as opposed to minimizing a cumulative cost over a trajectory.
@@
On the application side, our unified algorithm has broad applicability to a variety of problems, including constrained bandits, stream-based active learning, online hypothesis testing, and sequential $L_{1}$ model calibration, among others. In the following, we review existing results related to these representative examples. The details of how these problems can be incorporated into our framework will be presented in Section 4.
@@
Constrained bandits have been extensively studied in both academia and industry, with applications in areas such as online movie recommendation (Balakrishnan et al., ), news recommendation (Li et al., 2010), online pricing (Ao et al., 2025), and market equilibrium computation under incomplete information (Jalota et al., 2023). Various types of constrained stochastic bandits have been explored, depending on the nature of the constraints. For example, knapsack bandits (Badanidiyuru et al., 2018; Wu et al., 2015; Kumar and Kleinberg, 2022) model scenarios where pulling each arm yields both a reward and a random consumption of a global budget. The objective is to maximize cumulative rewards before the budget is exhausted. Another common formulation involves each arm being associated with both a reward distribution and a cost distribution, where the goal is to maximize cumulative rewards while ensuring that the expected cost at each round stays below a specified threshold—this is known as stage-wise constraints, which is also the focus of this paper. Several studies have addressed linear bandits with stage-wise constraints using various approaches, including explore-exploit (Amani et al., 2019), UCB (Pacchiano et al., 2021), and Thompson Sampling (Moradipari et al., 2021). In addition, Wang et al. (2022) investigated pure exploration under constraints in bandit settings. However, prior work has primarily focused on non-contextual problems. In contrast, we study decision-making with arbitrary context distributions and introduce a significantly more general framework that goes beyond optimizing for mean rewards alone.
@@
Online Hypothesis Testing concerns making sequential decisions between the null and alternative hypotheses given informative inputs at each round $t$, a problem originally studied by Wald and Wolfowitz (1948) and Robbins (1952). Since then, extensive research has focused on minimizing the expected experiment length for learning near-optimal decision rules (Naghshvar and Javidi, 2013; Chang and Bloch, 2023). Pan et al. (2022) investigated sequential hypothesis testing under type I and type II error constraints, while Lan and Wang (2021) introduced an active learning-style approach by incorporating budget constraints on the number of queries. In contrast to these works, which aim to design optimal policies under varying resource constraints, Malek et al. (2017) studied fixed-horizon multiple hypothesis testing with type I error control. On the application side, sequential hypothesis testing plays a critical role in clinical trials, where it is essential to maximize the probability of correctly diagnosing conditions while controlling the risk of missed diagnoses (Bartroff et al., 2012; Jennison and Turnbull, 1999). Our proposed algorithm is capable of handling sequential hypothesis testing under any fixed horizon setting. See  Section 4.3 for details.
@@
Sequential Model Calibration aims to ensure that a model’s predictions are unbiased, conditional on the predicted value. Roth (2024) provided a comprehensive overview of various calibration error metrics, including average $L_{1}$, $L_{2}$, and $L_{\infty}$ calibration errors, as well as quantile calibration errors (Gopalan et al., 2021; Úrsula Hébert-Johnson et al., 2018). The first sequential calibration algorithm was introduced by Foster and Vohra (1998). Later,  Foster (1999); Úrsula Hébert-Johnson et al. (2018) proposed the notion of multicalibration and developed corresponding algorithms. Additionally, Abernethy et al. (2011) established the equivalence between calibration and Blackwell approachability. Model calibration is also closely related to other areas of uncertainty quantification, such as conformal prediction (Bastani et al., 2022). Within our framework  Section 4.4, the proposed algorithm can be directly applied to sequential model calibration tasks with $L_{1}$ calibration error.
3 Problem Setup
3.1 General Online Decision Making with Constraints
@@
In this section, we formalize our constrained sequential decision-making problem. The decision maker (DM) operates over a context space $\mathcal{X}$ and a finite action space $\mathcal{A}$**We could extend our algorithm to infinite action spaces by invoking the counterfactual action divergences concept. We refer readers to Xu and Zeevi (2020) for details about action divergences. For each state-action pair $(x,a)$, there are two associated random variables with unknown distributions: $f^{*}_{x,a}\in\mathcal{F}$ and $g^{*}_{x,a}\in\mathcal{G}$. Here, $\mathcal{F}$ and $\mathcal{G}$ are known probability density model classes, which we refer to as the utility density class and the constraint density class, respectively.
@@
In general, a stochastic policy at time $t$ is a mapping $\pi_{t}:\mathcal{X}\rightarrow\Delta(\mathcal{A})$. However, since our sequential decision-making problem is per-context, at each round $t$, the DM only needs to specify the distribution $\pi_{t}(\cdot|x_{t})$ for the realized context $x_{t}$, rather than for the entire space $\mathcal{X}$. For notational simplicity, we will sometimes refer to $\pi_{t}(\cdot|x_{t})$ itself as the policy.
@@
where $\tau$ is some known threshold number.
However, the DM does not have access to the true distributions $f^{*}$ and $g^{*}$ and must instead estimate them from collected data. Due to the inherent stochasticity and estimation errors, it is generally impossible to guarantee feasibility with certainty. Therefore, in this paper, we focus on high-probability feasibility: the DM must ensure that, for some $\delta>0$, with probability at least $1-\delta$, the selected policy sequence $\left\{\pi_{t}\right\}_{t=1}^{T}$ is feasible for all $t\in[T]$. For any natural number $n$, we use $[n]$ to denote the number set $\left\{1,2,\cdots,n\right\}$.
@@
where $\pi^{*}$ denotes the optimal feasible stochastic policy, and $\pi_{t}$ is the policy selected by the DM at round $t$.
@@
Furthermore, to ensure that a feasible action always exists at each round $t$, we assume the presence of a safe action $a_{0}$ for every context $x\in\mathcal{X}$. Specifically, the DM has prior knowledge that this safe action satisfies $\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=c_{0}$ and $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})\leq\tau$ for all $x$. This guarantees that the DM can always select at least one policy that satisfies the constraint, regardless of the uncertainty in estimating $f^{*}$ and $g^{*}$. Such an assumption is standard in constrained decision-making settings to prevent infeasibility arising from distributional uncertainty  (Pacchiano et al., 2021; Kumar and Kleinberg, 2022).
@@
$\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=r_{0},\ \text{and}\ \tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})=c_{0}\leq\tau$

### v3

@@ Abstract
Contextual online decision-making problems with constraints appear in a wide range of real-world applications, such as adaptive experimental design under safety constraints, personalized recommendation with resource limits, and dynamic pricing under fairness requirements. In this paper, we investigate a general formulation of sequential decision-making with stage-wise feasibility constraints, where at each round, the learner must select an action based on observed context while ensuring that a problem-specific feasibility criterion is satisfied. We propose a unified algorithmic framework that captures many existing constrained learning problems, including constrained bandits, active learning with label budgets, online hypothesis testing with Type I error control, and model calibration. Central to our approach is the concept of upper counterfactual confidence bounds, which enables the design of practically efficient online algorithms with strong theoretical guarantees using any offline conditional density estimation oracle. To handle feasibility constraints in complex environments, we introduce a generalized notion of the eluder dimension, extending it from the classical setting based on square loss to a broader class of metric-like probability divergences. This allows us to capture the complexity of various density function classes and characterize the utility regret incurred due to feasibility constraint uncertainty. Our result offers a principled foundation for constrained sequential decision-making in both theory and practice.
@@ 1 Introduction
Contextual sequential decision-making plays a critical role in various domains, including online platforms (Ban et al., 2024), healthcare systems (Zhou et al., 2023), and other industries. In many real-world applications, the decision maker (DM) often operates under some constraints that must be satisfied (Min and Azizan, 2024). For example, there may be safety constraints in power systems (Ceusters et al., 2023), fairness constraints in online platforms (Verma et al., 2024), or risk constraints in portfolio management (Galichet et al., 2013). The presence of such constraints significantly complicates the sequential decision-making process, making it more challenging to manage the inherent exploration-exploitation trade-off.
To address this challenge, for different statistical tasks, a variety of adaptive constrained decision-making algorithms have been proposed, e.g., constrained bandits (Jagerman et al., 2020), active learning under constraints (Awasthi et al., 2021), and online hypothesis testing with Type I error control (Malek et al., 2017). Despite the diversity of these applications, they often share a common structural pattern. In each round, the DM observes a context $x_{t}$, such as the symptoms of the patient in a healthcare setting, and based on this information, the DM selects an action $a_{t}$ that is deemed feasible with respect to some estimated constraint. Subsequently, the DM receives feedback in the form of data sampled from the underlying distributions associated with the context-action pair $(x_{t},a_{t})$ and proceeds to the next round.
@@
In this paper, we introduce a unified framework for general constrained sequential decision-making problems and present an efficient algorithm to solve them. More specifically, our contributions can be summarized as follows:
  • A New General Framework: We propose a new unifed framework to handle general online decision-making problems with stage-wise constraints (Section 3). Our framework handles many sequential decision-making problems, such as constrained bandits, stream-based active learning, online hypothesis testing, and sequential model calibration as special cases (Section 4).
  • A Novel Conceptual Generalization of Eluder Dimension: Eluder dimension (Russo and Van Roy, 2013) is a complexity measure often used to characterize the difficulty of optimal exploration in bandits and the regret caused by conservative exploration in constrained online learning (Sridharan and Yoo, 2024). We extend the idea and concept of the eluder dimension (Russo and Van Roy, 2013) from the reward function class with squared error to any density model class with any metric-like probability divergence (Section 5). This new measure helps us balance aggressive and conservative exploration to handle feasibility.
  • An Efficient Unified Algorithm: We propose a unified algorithm GED-UCB that could easily transfer any offline density estimation oracle into an efficient sequential decision-making algorithm (Section 6). Unlike Hu et al. (2025), where the underlying density class is restricted to a linear model, our algorithm could handle any density model class, including linear and non-linear model classes. This greatly enriches the range of problems that we can solve.
@@ 2 Related Work
Sequential decision-making is a fundamental problem in modern machine learning. Early work, such as Bastani and Bayati (2020), focused on sequential decision-making with finite high-dimensional covariates, using contextual bandits as a representative example. From a more general perspective, Foster et al. (2021) and Zhong et al. (2022) investigated the learnability of general model classes by introducing the decision-estimation coefficient and its generalizations to characterize the complexity of interactive decision-making. Closely related, Hu et al. (2025) proposed an algorithm for unconstrained online decision-making problems based on linear functional regression. However, these studies focus primarily on maximizing expected rewards without addressing additional structural constraints. In contrast, our paper provides a more general framework that accommodates sequential decision-making under stage-wise constraints. Additionally, in contrast with Hu et al. (2025), we support both linear and nonlinear density classes while incorporating constraints into the decision process.
@@
The study of online learning with constraints was initiated by Mannor et al. (2009), where the learner aims to maximize expected rewards subject to long-term sample path constraints. Later on, Mahdavi et al. (2012); Castiglioni et al. (2022); Liakopoulos et al. (2019); Yu et al. (2017) studied the regret upper bound and constraint violation in online optimization with long-term constraints. Instead of quantifying constraint violation, our paper gives a feasibility guarantee such that with high probability, it holds for all rounds. On the other hand, Sridharan and Yoo (2024); Hutchinson et al. (2024) investigated online optimization with per-round constraints. The benchmark in these papers is the best fixed action in hindsight, which is non-contextual and the regret is static, whereas in our paper, we study the per-context optimal decision-making problem and our benchmark is contextualized dynamic regret. In brief, we have a high probability feasibility guarantee with a dynamic regret guarantee in our paper.
@@
Our work is also closely related to the body of research on online learning with general function approximation. Xu and Zeevi (2020) proposed the Upper Counterfactual Confidence Bound (UCCB) algorithm, which constructs upper confidence bounds in the policy space under general reward function approximation. Levy et al. (2024) used UCCB principle to study the regret in contextual MDPs. Foster and Rakhlin (2020) addressed the contextual bandit problem using inverse-gap weighting policies, assuming access to an online regression oracle. More recently, Pacchiano (2024) and Jia et al. (2024) examined general contextual bandits with small noise levels and established connections between reward variance and regret. In the context of reinforcement learning, Wang et al. (2020a, b) studied regret minimization under general value function approximation. In contrast, our work focuses on general sequential decision-making with feasibility constraints, leveraging density function approximation oracles.
@@
Another related literature is safe reinforcement learning (RL). For comprehensive overviews, we refer readers to the surveys by Gu et al. (2022) and Wachi et al. (2024). Wachi and Sui (2020) investigated reinforcement learning with explicit state-wise safety constraints, and Li and Azizan (2024) studied its multi-agent variant and convergence to Nash equilibrium. Furthermore, HasanzadeZonuzy et al. (2021); Amani et al. (2021a, b) analyzed the sample complexity and regret of learning in value function constrained Markov Decision Processes (MDPs). Additionally, Miryoosefi et al. (2019) studied constrained MDPs where policies are required to lie within a convex policy set. Although contextual bandits could technically be viewed as a special case of RL by defining the state to be the context and setting the horizon to be $1$, the exploration of RL algorithms could be unnecessarily inefficient in this setting and their statistical guarantees typically scale with the cardinality of the state space, whereas in our results, there is no dependence on context space in the regret bound. Additionally, our framework accommodates a broader class of feasibility constraints than value-based ones.
@@
Furthermore, we remark on some connections and differences between our paper and control theory. Regret-optimal control (Hazan et al., 2020; Sabag et al., 2023; Goel and Hassibi, 2021b) is about minimizing the total cost relative to a clairvoyant optimal policy that has foreknowledge of the system dynamics. Recently, Martin et al. (2022) studied this problem under a safety constraint. In the literature on regret-optimal control, the primary focus has been on identifying the optimal controller and characterizing its structural properties (Sabag et al., 2021; Goel and Hassibi, 2021a; Martin et al., 2024). Moreover, optimal control focuses on minimizing cumulative costs generated by a dynamic system along a single trajectory (Li et al., 2019). In our framework, however, contexts do not have system dynamics; so, we aim to maximize cumulative rewards through repeated online experimentation, as opposed to minimizing a cumulative cost over a trajectory.
2.1 Special Cases
@@
As mentioned earlier, our unified algorithm has broad applicability to a variety of problems, including constrained bandits, stream-based active learning, online hypothesis testing, and sequential $L_{1}$ model calibration, among others. In the following, we review existing results related to these representative examples. The details of how these problems can be viewed as special cases of our framework will be presented in Section 4.
@@
Constrained bandits have been extensively studied from both theoretical and applied perspectives, with applications in areas such as online movie recommendation (Balakrishnan et al., 2018), news recommendation (Li et al., 2010), online pricing (Ao et al., 2025), and market equilibrium computation under incomplete information (Jalota et al., 2023). Various types of constrained stochastic bandits have been explored, depending on the nature of the constraints. For example, knapsack bandits (Badanidiyuru et al., 2018; Wu et al., 2015; Kumar and Kleinberg, 2022) model scenarios where pulling each arm yields both a reward and a random consumption of a global budget. The objective is to maximize cumulative rewards before the budget is exhausted. Another common formulation involves each arm being associated with both a reward distribution and a cost distribution, where the goal is to maximize cumulative rewards while ensuring that the expected cost at each round stays below a specified threshold—this is known as stage-wise constraints, which is also the focus of this paper. Several studies have addressed linear bandits with stage-wise constraints using various approaches, including explore-exploit (Amani et al., 2019), UCB (Pacchiano et al., 2021), and Thompson Sampling (Moradipari et al., 2021). In addition, Wang et al. (2022) investigated pure exploration under constraints in bandit settings. However, prior work has primarily focused on non-contextual problems. In contrast, we study decision-making with arbitrary context distributions and introduce a significantly more general framework that goes beyond optimizing for mean rewards alone (see Section 4.1).
@@
Online Hypothesis Testing concerns making sequential decisions between the null and alternative hypotheses given informative inputs at each round, a problem originally studied by Wald and Wolfowitz (1948) and Robbins (1952). Since then, extensive research has focused on minimizing the expected experiment length for learning near-optimal decision rules (Naghshvar and Javidi, 2013; Chang and Bloch, 2023). Pan et al. (2022) investigated sequential hypothesis testing under type I and type II error constraints, while Lan and Wang (2021) introduced an active learning-style approach by incorporating budget constraints on the number of queries. In contrast to these works, which aim to design optimal policies under varying resource constraints, Malek et al. (2017) studied fixed-horizon multiple hypothesis testing with type I error control. On the application side, sequential hypothesis testing plays a critical role in clinical trials, where it is essential to maximize the probability of correctly diagnosing conditions while controlling the risk of missed diagnoses (Bartroff et al., 2012; Jennison and Turnbull, 1999). Our proposed algorithm is capable of handling sequential hypothesis testing under any fixed horizon setting (see Section 4.3 for details).
@@
Sequential Model Calibration aims to ensure that a model’s predictions are unbiased, conditional on the predicted value. Roth (2024) provided a comprehensive overview of various calibration error metrics, including average $L_{1}$, $L_{2}$, and $L_{\infty}$ calibration errors, as well as quantile calibration errors (Gopalan et al., 2021; Úrsula Hébert-Johnson et al., 2018). The first sequential calibration algorithm was introduced by Foster and Vohra (1998). Later, Foster (1999); Úrsula Hébert-Johnson et al. (2018) proposed the notion of multicalibration and developed corresponding algorithms. Additionally, Abernethy et al. (2011) established the equivalence between calibration and Blackwell approachability. Model calibration is also closely related to other areas of uncertainty quantification, such as conformal prediction (Bastani et al., 2022). Within our framework, the proposed algorithm can be directly applied to sequential model calibration tasks with $L_{1}$ calibration error (see Section 4.4).
3 General Online Decision Making with Constraints
@@
In this section, we formalize our constrained sequential decision-making problem. The decision maker (DM) operates over a context space $\mathcal{X}$ and a finite action space $\mathcal{A}$**We could extend our algorithm to infinite action spaces by invoking the counterfactual action divergences concept. We refer readers to Xu and Zeevi (2020) for details about action divergences.. For each state-action pair $(x,a)$, there are two associated random variables with unknown distributions $f^{*}_{x,a}\in\mathcal{F}$ and $g^{*}_{x,a}\in\mathcal{G}$. Here, $\mathcal{F}$ and $\mathcal{G}$ are known probability density model classes, which we refer to as the utility density class and the constraint density class, respectively.
@@
In general, a stochastic policy at time $t$ is a mapping $\pi_{t}:\mathcal{X}\rightarrow\Delta(\mathcal{A})$. However, since our sequential decision-making problem is per-context, at each round $t$, the DM only needs to specify the distribution $\pi_{t}(\cdot|x_{t})$ for the realized context $x_{t}$, rather than for the entire space $\mathcal{X}$. For simplicity, we will sometimes refer to $\pi_{t}(\cdot|x_{t})$ itself as the policy.
@@
where $\tau$ is some known threshold value.
However, the DM does not have access to the true distributions $f^{*}$ and $g^{*}$ and must instead estimate them from collected data. Due to the inherent stochasticity and estimation errors, it is generally impossible to guarantee feasibility with certainty. Therefore, in this paper, we focus on high-probability feasibility: the DM must ensure that, for any prespecified $\delta>0$, with probability at least $1-\delta$, the selected policy sequence $\left\{\pi_{t}\right\}_{t=1}^{T}$ is feasible for all $t\in[T]$. We use $[n]$ to denote the set $\left\{1,2,\cdots,n\right\}$ for any natural number $n$.
@@
where $\pi^{*}$ denotes the optimal feasible stochastic policy, and $\pi_{t}$ is the policy selected by the DM at round $t$. By saying a policy $\pi$ is feasible at round $t$ after receiving context $x_{t}$, we mean that $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x_{t},\pi})\leq\tau$. As a remark, we use $\mathbb{E}_{X\sim\mathbb{P}}$ to denote the expectation taken with respect to the random variable $X$ with distribution $\mathbb{P}$.
@@
Furthermore, to ensure that a feasible action always exists at each round $t$, we assume the presence of a safe action $a_{0}$ for every context $x\in\mathcal{X}$. Specifically, the DM has prior knowledge that this safe action satisfies $\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=c_{0}$ and $\tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})\leq\tau$ for all $x$. This guarantees that the DM can always select at least one policy that satisfies the constraint. Such an assumption is standard in constrained decision-making settings to prevent infeasibility arising from distributional uncertainty (Pacchiano et al., 2021; Kumar and Kleinberg, 2022).
@@
$\tilde{{\mathcal{T}}_{1}}(f^{*}_{x,a_{0}})=r_{0},\ \text{and}\ \tilde{{\mathcal{T}}_{2}}(g^{*}_{x,a_{0}})=c_{0}\leq\tau.$